# Fast Diffusion-Based Counterfactuals for Shortcut Removal and Generation

### v1

@@ Abstract
Shortcut learning is when a model – e.g. a cardiac disease classifier – exploits correlations between the target label and a spurious shortcut feature, e.g. a pacemaker, to predict the target label based on the shortcut rather than real discriminative features. This is common in medical imaging, where treatment and clinical annotations correlate with disease labels, making them easy shortcuts to predict disease. We propose a novel detection and quantification of the impact of potential shortcut features via a fast diffusion-based counterfactual image generation that can synthetically remove or add shortcuts. Via a novel inpainting-based modification we spatially limit the changes made with no extra inference step, encouraging the removal of spatially constrained shortcut features while ensuring that the shortcut-free counterfactuals preserve their remaining image features to a high degree. Using these, we assess how shortcut features influence model predictions.
This is enabled by our second contribution: An efficient diffusion-based counterfactual explanation method with significant inference speed-up at comparable image quality as state-of-the-art. We confirm this on two large chest X-ray datasets, a skin lesion dataset, and CelebA.
**footnotetext: Equal contribution
@@ 1 Introduction
Figure 1: An illustration of proposed visual counterfactual generation and proposed shortcut learning detection pipeline.
Shortcut learning [24] denotes the situation in which a model exploits a spurious correlation between a ‘shortcut’ feature $s$ and the target outcome $y$. Known examples include grass being used as a shortcut for sheep, or surgical skin markings being shortcuts for malignant skin lesions [71]. Since the correlation is only spurious, and not causal, shortcut learning results in misleadingly high model performance on test data that also contains shortcuts, but poor ability to generalize to data without shortcuts. As flexible models can easily and accurately identify scores of potential (often non-obvious) shortcut features from images [48, 31, 25], shortcut learning is a key obstacle on the path to image classification models that generalize well. This is especially true in medical imaging, where classification tasks often suffer from high intrinsic uncertainty and shortcuts are highly effective in improving classification performance [71, 49, 5, 31, 18, 35].
Detecting shortcut learning is non-trivial. Stratified analyses of model performance on samples with and without the potential shortcut might yield some indication, but cannot confirm shortcut learning as any observed performance differences could be caused by other distributional differences. Explainable AI (XAI) methods have been proposed to obtain more concrete evidence [72, 65, 66, 51], but require manual inspection of explanations making their use further complicated by explanation instability [3, 64] and infinite sets of feasible competing explanations [53, 40].
We propose a novel approach leveraging diffusion-based counterfactual image generation [33, 4] to generate ‘shortcut counterfactuals’, that do or do not contain a shortcut feature of interest. Different from standard counterfactual explanations which change the target label, we rather seek to change the shortcut label. By assessing how predictions change with addition or removal of the shortcut feature, we quantify the model’s degree of shortcut learning (see Fig. 1).
@@
    A fast diffusion-based method FastDiME for visual counterfactual explanations using denoised images instead of generated ones to obtain the gradients during sampling. This enables significant inference speed-up of up to 20 times faster than current methods, with comparable counterfactual quality.
@@
    Leveraging a novel mask-based inpainting scheme during sampling to limit changes between the original and the counterfactual image to a small region, we prevent the removal of discriminative features.
@@
    A novel pipeline for detecting and quantifying shortcut learning via the generation of shortcut counterfactuals from our diffusion-based counterfactual method.
@@
    A demonstration of the generality of our method for counterfactual explanations as well as its utility in detecting and quantifying shortcut learning in multiple realistic scenarios with datasets from different domains.
@@ 2 Related work
Figure 2: Proposed FastDiME method. In each step, noised image $x_{t}^{c}$ is sampled with the guidance of the counterfactual loss, leveraging information derived from the denoised image $\bar{x}_{t}^{c}$. A mask is applied to prevent changes in regions less relevant to the task.
@@ 3 Methods
Counterfactual image explanation aims to solve the following problem: Assume given an image classification problem with a particular class $c$ of interest – for instance, whether or not a chest X-ray contains a cardiac pacemaker – and an image $x$. Can we provide an updated image $x^{c}$ which remains as close as possible to $x$ while both being visually realistic and undergoing sufficient visual change to clearly belong to the class $c$? In our medical imaging case, this could consist of artificially adding or removing the pacemaker without changing any of the remaining patient anatomy.
@@ 3.1 Diffusion Models for Counterfactual Explanations (DiME)
Jeanneret et al. [33] utilize diffusion models for generating counterfactual explanations through a guided diffusion process. The image is guided towards the desired counterfactual class using the classifier’s loss objective $L_{c}$, while the counterfactual $x^{c}$ is constrained to remain close to the original $x$ through an $L_{1}$ loss and a perceptual loss $L_{perc}$. The overall gradient for the counterfactual loss term thus is of the form $\triangledown_{CF}=\lambda_{c}\triangledown L_{c}+\lambda_{1}\triangledown L_{1}+\lambda_{p}\triangledown L_{perc}$, where $\lambda_{c}$, $\lambda_{1}$ and $\lambda_{p}$ are hyperparameters, and $L_{1}$ is measured between a noisy version and the original image. To obtain meaningful gradients with their classifier, another nested diffusion process is used to synthesize (unconditionally) an image at the cost of running a whole DDPM process for each step. DiME  [33] is summarized as follows:
  • Corrupt input up to noise level $\tau$ with the forward process
@@
      Compute the gradients $\triangledown_{CF}(\hat{x}^{c}_{t})$ based on a clean image $\hat{x}^{c}_{t}$ synthesized using an inner forward process continuing the unconditional generation process from the current step $t$, i.e, $\hat{x}^{c}_{t}=\text{DDPM(}x^{c}_{t},t)$
@@ 3.2 Fast generation of high-quality counterfactuals
We now present FastDiME, which improves on DiME using an efficient gradient estimation and a novel mask-based inpainting scheme. An illustration can be found in Fig. 2.
@@
As explained above, gradients in the original DiME model are retrieved from generated images by re-running the entire DDPM process. This is computationally expensive, with complexity $\mathcal{O}(T^{2})$. To speed up inference while maintaining image quality, we propose to use the denoised $\bar{x}^{c}_{t}$ for calculating the gradients. In the forward diffusion process ($q$), the noisy image $x_{t}$ from $x_{1}$ is synthesized using a Gaussian distribution at each timestep with scheduled variance $\beta$:
@@
As derived in Ho et al. [28], a direct noising process conditioned on the input $x^{c}_{0}$ is feasible by marginalizing Eq. (1),
@@
$x^{c}_{t}=\bar{\alpha}_{t}x^{c}_{0}+\sqrt{(1-\bar{\alpha}_{t})}\epsilon,$ (3)
where $\alpha_{t}\coloneqq 1-\beta_{t}$, $\bar{\alpha}_{t}\coloneqq\prod_{s=0}^{t}\alpha_{s}$ and $\epsilon\sim\mathcal{N}(0,\boldsymbol{I})$. Given a noise estimate $\bar{\epsilon}$ (from a denoiser), we can use Eq. (3) to retrieve the estimated $\bar{x}^{c}_{0}$ from the noisy image $x^{c}_{t}$:
@@
$\bar{x}^{c}=\bar{x}^{c}_{0}=\mathop{\mathbb{E}}[\hat{x}^{c}_{0}],$
which we use to calculate the gradients in our counterfactual generation steps. The time complexity of adopting a denoised image is $\mathcal{O}(T)$, which significantly expedites the entire inference procedure. Further experiments on the time consumption can be found in  Sec. 4.3.
@@
Refer to caption Refer to caption Refer to caption
Figure 3: Toy 2D example for visualization of the counterfactual generation convergence comparing DiME and FastDiME. We place two attributes (A, B) for our shortcut. Both techniques are initialized with $x_{t}$ (marked in red) and the target class is the distribution centered at attribute A. Blue vectors indicate the gradients of classification loss $\triangledown L_{c}$ at each step. When calculating this gradient at each step, DiME uses a new unconditional sample from DDPM, which could lead to noisy gradients. In contrast, FastDiME uses the expected image $\bar{x}_{0}$ at each step to calculate the gradients, which results to more stable and faster convergence.
Mask-based inpainting.
Our application of counterfactual generation is to remove shortcut features from images. Shortcut features tend to be highly localized; thus, it is often desirable to constrain changes to the image made by the counterfactual generation process to a small region of the image. To achieve this goal, Jeanneret et al. [34] performs an extra inpainting step after the usual counterfactual generation step using RePaint [43] to ensure highly localized changes to the image. However, due to the in-painting step lacks classifier guidance, their approach does not guarantee that the generated images remain valid counterfactuals – in principle, the class label could change again during the inpainting step. Inspired by their approach, we integrate in-painting directly into the initial counterfactual generation process to preserve more features from the original image, while incorporating other terms in the optimization. The inpainting is performed by masking the sampled image and denoised images at time-step $t$, i.e.,
@@
where $x_{t}\sim q(x_{t}|x_{0},t)$ refers to the sample from the forward process on the original image $x_{0}$ and adds a Gaussian noise to it. Different from [34], our inpainting approach is fully gradient-guided and ensures the validity of the counterfactuals (see Appendix C). We implement this inside our main process after a warm-up period of $\tau_{w}$ time-steps, with $1\leq\tau_{w}\leq\tau$. We further experiment with 2-step approaches keeping $M$ fixed after a completed run of guided diffusion with efficient gradient estimations or a completed run of our full method including our inpainting scheme. We refer to these as FastDiME-2 and FastDiME-2+, respectively.
@@ 3.3 Shortcut learning detection
Figure 4: Proposed shortcut learning detection pipeline. We construct three synthetic training datasets with varying degrees of correlation between the shortcut feature $s$ and the target label $y$. If model predictions differ strongly between natural images and shortcut counterfactual images (while leaving the target label unchanged), shortcut learning has occurred.
Assume a classifier $f(x)=\hat{y}$ that we suspect of shortcut learning, different from the one used for counterfactual image generation. For example, $f$ trained to diagnose lung diseases from chest X-ray images but instead might use medical equipment such as cardiac pacemakers or chest drains as shortcuts. We propose the following pipeline for quantifying the degree to which $f$ has indeed learned to rely on the shortcut feature (see Fig. 4):
@@
    Curate natural test sets, along with a test set where the suspected shortcut feature has been counterfactually flipped using our method FastDiME. In particular, curate
@@ 4.1 Datasets and implementation details
For CelebA, we use the same trained DenseNet121 classifier [29] and DDPM [28] as in DiME [33] and ACE [34] for fair comparisons.
For all variants of our method, namely, FastDiME, FastDiME-2, and FastDiME-2+, we follow the same hyperparameters as in DiME [33]. For the medical datasets, we train DDPMs with 1000 steps, use UNet’s [59] encoder architecture as the shortcut classifier for counterfactual generation, and set $\tau$ to 160 out of 400 re-spaced time steps without including $L_{perc}$ in the counterfactual loss during inference. For all datasets, we set $\tau_{w}=\frac{\tau}{2}$, and normalize, threshold and dilate the masks similar to ACE [34]. For the shortcut detection experiments, the task classifier suspected of shortcut learning is a ResNet-18 [26]. Our source code for all experiments will be publicly available.

### v2

@@ Abstract
Shortcut learning is when a model – e.g. a cardiac disease classifier – exploits correlations between the target label and a spurious shortcut feature, e.g. a pacemaker, to predict the target label based on the shortcut rather than real discriminative features. This is common in medical imaging, where treatment and clinical annotations correlate with disease labels, making them easy shortcuts to predict disease. We propose a novel detection and quantification of the impact of potential shortcut features via a fast diffusion-based counterfactual image generation that can synthetically remove or add shortcuts. Via a novel self-optimized masking scheme we spatially limit the changes made with no extra inference step, encouraging the removal of spatially constrained shortcut features while ensuring that the shortcut-free counterfactuals preserve their remaining image features to a high degree. Using these, we assess how shortcut features influence model predictions.
This is enabled by our second contribution: An efficient diffusion-based counterfactual explanation method with significant inference speed-up at comparable image quality as state-of-the-art. We confirm this on two large chest X-ray datasets, a skin lesion dataset, and CelebA. Our code is publicly available at fastdime.compute.dtu.dk.
Keywords:
Shortcut Learning Counterfactuals Diffusion Models
@@ 1 Introduction
footnotetext: N. Weng and P. Pegios contributed equally to this work.
Figure 1: Shortcut detection: SmoothGrad and CF explanation (Left), as two XAI methods, indicate which region in the image could influence the model decision (e.g. from disease to non-disease). Although this includes the shortcut features, it does not clearly indicate it. Therefore an expert is required for further visual inspection. Our counterfactual approach (Right) only removes the desired shortcut attribute. With this we can validate if that specific attribute played a role in the model decision. SmoothGrad visualization: highlighting areas crucial to the model decisions. CF explanation: difference map of the original image and CF (blue/red: information removal/addition).
Shortcut learning [25] denotes the situation in which a model exploits a spurious correlation between a ‘shortcut’ feature $s$ and the target outcome $y$. Known examples include grass being used as a shortcut for sheep, or surgical skin markers being shortcuts for malignant skin lesions [75]. Since the correlation is only spurious, and not causal, shortcut learning results in misleadingly high model performance on validation data that also contains shortcuts, but poor ability to generalize to data without shortcuts. As prediction targets are often more complex than shortcuts, deep models can more easily and accurately identify the shortcuts in the data [51, 33, 27]. Therefore, shortcut learning is a key obstacle to achieving well-generalized image classification models. This is especially true in medical imaging, where classification tasks often suffer from high intrinsic uncertainty and shortcuts are highly effective in improving classification performance [75, 52, 5, 33, 19, 37].
Detecting shortcut learning is non-trivial. Stratified analyses of model performance on samples with and without the potential shortcut might yield some indication, but cannot confirm shortcut learning as any performance gaps could result from other distributional differences. Explainable AI (XAI) methods have been proposed to obtain more concrete evidence [76, 69, 70, 54], but require manual inspection of explanations, complicating further use by explanation instability [3, 68] and infinite sets of feasible competing explanations [56, 43] (see Fig. 1).
We propose a novel approach leveraging diffusion-based counterfactual (CF) generation [35, 4] to generate ‘shortcut counterfactuals’, that do or do not contain a shortcut feature of interest. Different from standard counterfactual explanations which change the target label, we rather seek to change the shortcut label. By assessing how predictions change with addition or removal of the shortcut feature, we quantify the model’s degree of shortcut learning (see Fig. 1).
@@
    A diffusion-based method FastDiME for counterfactual generation, employing approximate gradients in sampling, achieving a 20x speedup over comparable state-of-the-art models while maintaining counterfactual quality.
@@
    A novel self-optimized masking scheme that confines counterfactual changes to a small region, with which we mitigate the unintentional removal of non-shortcut features in counterfactual images.
@@
    A novel pipeline for detecting and quantifying shortcut learning via the generation of shortcut counterfactuals from our diffusion-based method.
@@
    A demonstration of the generality of our method for counterfactual generation as well as its utility in detecting and quantifying shortcut learning in multiple realistic scenarios from different domains.
@@ 2 Related work
Guided diffusion and image editing.
Many approaches have been built to manipulate input images. From these techniques, the most relevant works include Universal Guidance [7], Motion Guidance [26], and GMD[39]. These methods improve the efficiency of the DiME approach by feeding the denoised image into the guiding classifier, and backpropagating the classifier gradients through the denoiser to calculate the gradients wrt. the input image. We, instead, use the denoised image to approximate the gradients for the input image and avoid a heavy back propagation step. The denoised image gradients act as a surrogate for the image gradients as they asymptotically reach the same image.
@@ 3 Methods
Counterfactual image explanation aims to solve the following problem: Assume given an image classification problem with a particular class $c$ of interest – for instance, whether or not a chest X-ray contains a cardiac pacemaker – and an image $x$ that does not belong to class $c$. Can we provide an updated image $x^{c}$ that remains as close as possible to $x$ while both being visually realistic and undergoing sufficient visual change to clearly belong to the class $c$? In our medical imaging case, this could consist of artificially adding or removing the pacemaker without changing any of the remaining patient anatomy.
@@ 3.1 Diffusion Models for Counterfactual Explanations (DiME)
Jeanneret et al. [35] utilize DDPMs for generating counterfactual explanations through a guided diffusion process [21]. The image is guided towards the counterfactual class using the classifier’s loss objective $L_{c}$, while the counterfactual $x^{c}$ is constrained to remain close to the original $x$ through an $L_{1}$ loss and a perceptual loss $L_{perc}$. The overall gradient for the counterfactual loss term is of the form $\triangledown_{CF}=\lambda_{c}\triangledown L_{c}+\lambda_{1}\triangledown L_{1}+\lambda_{p}\triangledown L_{perc}$, where $\lambda_{c}$, $\lambda_{1}$ and $\lambda_{p}$ are hyperparameters, and $L_{1}$ is measured between a noisy version and the original image. To obtain meaningful gradients with their classifier, another nested diffusion process is used to synthesize (unconditionally) an image at the cost of running a whole DDPM process for each step. DiME [35] is summarized as follows:
  • Corrupt input up to noise level $\tau$ with the forward process.
@@
      Compute the gradients $\triangledown_{CF}(\hat{x}^{c}_{t})$ based on a clean image $\hat{x}^{c}_{t}$ synthesized using an inner forward process continuing the unconditional generation process from the current step $t$, i.e, $\hat{x}^{c}_{t}=\text{DDPM(}x^{c}_{t},t)$.
@@
Figure 2: Proposed FastDiME method. In each step, noised image $x_{t}^{c}$ is sampled with the guidance of the counterfactual loss, leveraging information derived from the denoised image $\bar{x}_{t}^{c}$. A self-optimized mask is automatically extracted and applied to prevent changes in regions less relevant to the task at each time step.
@@ 3.2 Fast generation of high-quality counterfactuals
We present FastDiME, which improves on DiME with an efficient gradient estimation and a novel self-optimized masking scheme (see Fig. 2).
@@
As explained above, gradients in the original DiME model are retrieved from generated images by re-running the entire DDPM process. This is computationally expensive, with complexity $\mathcal{O}(T^{2})$. To speed up inference while maintaining image quality, we propose using the denoised $\bar{x}^{c}_{t}$ to calculate the gradients. In the forward diffusion process ($q$), the noisy image $x_{t}$ from $x_{t-1}$ is synthesized using a Gaussian distribution at each timestep with scheduled variance $\beta$:
@@
As derived by Ho et al. [30], a direct noising process conditioned on the input $x^{c}_{0}$ is feasible by marginalizing Eq. (1),
@@
$x^{c}_{t}=\sqrt{\bar{\alpha}_{t}}x^{c}_{0}+\sqrt{(1-\bar{\alpha}_{t})}\epsilon,$ (3)
where $\alpha_{t}\coloneqq 1-\beta_{t}$, $\bar{\alpha}_{t}\coloneqq\prod_{s=0}^{t}\alpha_{s}$ and $\epsilon\sim\mathcal{N}(0,\boldsymbol{I})$. Given a noise estimate $\bar{\epsilon}$ (from a denoiser), we can use Eq. (3) to retrieve the estimated denoised $\bar{x}^{c}_{t}$ at time-step $t$ from the noisy image $x^{c}_{t}$:
@@
$\bar{x}^{c}=\bar{x}^{c}_{0}=\mathop{\mathbb{E}}[\hat{x}^{c}_{0}],$ (5)
which we use to calculate the gradients in counterfactual generation steps. The gradients from $\bar{x}^{c}_{t}$ act as a surrogate for the image gradients $x^{c}_{t}$ as both images asymptotically reach the same solution in the diffusion process. The time complexity of adopting a denoised image is $\mathcal{O}(T)$, which significantly expedites the entire inference procedure (refer to Sec. 4.3 for time consumption experiments).
@@
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 3: Toy 2D example for visualization of the counterfactual generation convergence comparing DiME, GMD, and FastDiME. Given the two class distributions A and B, each method tries to bring the initial point $x_{t}$ (red point) to class A. Blue vectors indicate the gradients of classification loss $\triangledown L_{c}$ at each step. When calculating this gradient at each step, DiME uses a new unconditional sample from DDPM, which could lead to noisy gradients. In contrast, FastDiME uses the expected image $\bar{x}_{0}$ at each step to calculate the gradients, which results in a more stable convergence. The plot on the right shows the convergence of each method in terms of the distance to A. This is averaged over 100 runs and it indicates that FastDiME is accurate and significantly faster than both the other methods.
Self-optimized masking.
Our application of counterfactual generation is to remove shortcut features from images. Shortcut features tend to be highly localized; thus, it is often desirable to constrain changes to the image made by the counterfactual generation process to a small region of the image. To achieve this goal, Jeanneret et al. [36] perform an extra inpainting step after the usual counterfactual generation step using RePaint [46] to ensure highly localized changes to the image. However, due to the in-painting step lacking classifier guidance, their approach does not guarantee that the generated images remain valid counterfactuals – in principle, the class label could change again during the inpainting step. Inspired by their approach, we integrate self-optimized masking directly into the initial counterfactual generation process to preserve more features from the original image, while incorporating other terms in the optimization. To this end, we first extract a mask $M_{t}$ by binarizing the difference between denoised image at time-step $t$ and original input $x_{0}$, and then masking the sampled image $x^{c}_{t}$ and denoised images $\bar{x}^{c}_{t}$ accordingly, i.e.,
$M_{t}$ $=\delta(\bar{x}^{c}_{t},x_{0})$ (6)
@@
where $x_{t}\sim q(x_{t}|x_{0},t)$ refers to the sample from the forward process on the original image $x_{0}$ that adds a Gaussian noise to it, and $\delta$ represents the function used for extracting mask. Note that the mask is not given, but automatically generated by our pipeline, and it can vary with each time-step $t$. Different from [36], our fully gradient-guided masking approach ensures the validity of the counterfactuals. We implement this inside our main process after a warm-up period of $\tau_{w}$ time-steps, with $1\leq\tau_{w}\leq\tau$. We further experiment with 2-step approaches keeping $M$ fixed after a completed run of guided diffusion with efficient gradient estimations or a completed run of our full method including our self-optimized mask scheme. We refer to these as FastDiME-2 and FastDiME-2+, respectively.
@@ 3.3 Shortcut learning detection
Figure 4: Validating shortcut learning detection pipeline. In order to test the shortcut detecting ability by shortcut-counterfactuals, we construct three synthetic training datasets with varying degrees of correlation between the shortcut feature $s$ and the target label $y$ and train classifiers based on them (Left and Middle). By measuring the difference in confidence leveling between the original image and shortcut-counterfactual, the degree of shortcut learning is examined (Right). If model predictions differ strongly between natural images and shortcut counterfactual images (while leaving the target label unchanged), shortcut learning has occurred.
Assume a classifier $f(x)=\hat{y}$ that we suspect of shortcut learning, different from the one used for counterfactual image generation. For example, $f$ trained to diagnose lung diseases from chest X-ray images but instead might use medical equipment such as cardiac pacemakers or chest drains as shortcuts. We propose the pipeline for quantifying the degree to which $f$ has indeed learned to rely on the shortcut feature as shown in Fig. 4 (Right), where we generate the shortcut counterfactuals of all images, i.e. images that remove the pacemaker, and evaluate the alteration of predictions.
In order to validate the proposed counterfactual-based shortcut detection pipeline, we propose the following framework (see Fig. 4):
@@
    Curate natural test sets, along with a test set where the suspected shortcut feature is counterfactually flipped using FastDiME. In particular, curate
@@ 4.1 Datasets and implementation details
For CelebA, we use the same trained DenseNet121 classifier [31] and DDPM [30] as in DiME [35] and ACE [36] for fair comparisons. For all variants of our method, namely, FastDiME, FastDiME-2, and FastDiME-2+, we follow the same hyperparameters as in DiME [35]. For the medical datasets, we train DDPMs with 1000 steps, use UNet’s [62] encoder architecture as the shortcut classifier for counterfactual generation, set $\tau$ to 160 out of 400 re-spaced time steps, and compute $L_{1}$ between the denoised and original input without including $L_{perc}$. For all datasets, we set $\tau_{w}=\frac{\tau}{2}$, and normalize, threshold and dilate the masks similar to ACE [36]. For the shortcut detection experiments, the task classifier suspected of shortcut learning is a ResNet-18 [28].
4.2 Counterfactual explanations
Evaluation criteria.
Table 1: Results for CelebA. Results for DiVE [60], DiVE100, STEEX [34], DiME [35] and ACE [36] are from [36]. We compute BKL, MAD and FR metrics for diffusion-based methods, and highlight the best and second-best performances.